# Multi-CATE: Multi-Accurate Conditional Average Treatment Effect Estimation Robust to Unknown Covariate Shifts

### v1

@@ Results
Train Shift CForest S-learner DR-learner T-learner CForest S-learner DR-l T-learner
size degree OS OS OS Ridge Tree OS Ridge Tree RCT wRCT RCT wRCT RCT RCT wRCT
0 37.43 47.2 94.84 100.75 76.41 86.4 30.54 33.85 7.71 7.76 24.31 24.37 11.11 21.63 11.12
0.25 37.22 47.11 95.35 75.57 81.38 87.48 31.09 31.7 9.33 11.57 26.49 26.4 11.53 19.14 11.52
0.5 38.09 47.77 93.25 70.52 66.67 84.78 35.26 35.56 17.13 34.55 33.46 35.15 18 24.62 17.87
0.75 35.07 44.28 93.6 91.21 64.14 78.57 30.76 29.05 27.5 52.74 34.61 37.01 25.26 32.78 26.67
500 1 38.69 49.32 102.71 95.48 166.78 91.69 37.94 31.16 47.8 84.65 41.38 43.62 40.63 40.01 43.95
1.25 34.34 43.26 92.51 77.29 69.69 81.17 36.15 26.93 67.26 158.72 47.33 42.52 55.63 49.43 52.96
1.5 34.78 43.29 117.41 64.09 82.92 84.48 34.91 24.9 88.28 141.6 54.21 43.13 65.94 70.72 55.18
1.75 36.48 44.88 86.79 84.26 77.4 77.93 36.47 25.58 104.41 134.25 66.15 46.86 85.97 89.8 78.45
2 40.22 51.41 117.57 103.47 108.49 91.47 35.91 27.33 117.4 166.51 73.7 51.99 96.27 91.76 85.72
0 24.77 39.92 69.39 38.64 44.04 60.21 28.88 27.3 8.2 8.49 25.29 25.55 11.9 20.17 12.31
0.25 22.66 36.62 61.15 46.68 40.69 56.5 21.7 21.79 9.11 17.4 25.27 25.67 12.27 19.06 13.67
0.5 23.36 40.53 63.01 43.82 57.06 60.26 24.92 20.93 16.28 45.42 31.93 35.5 17.8 25.63 19.64
0.75 24.91 41.14 62.5 37.85 47.54 61.23 28.29 20.71 29.3 72.08 35.42 39.42 27.35 38.08 30.23
2000 1 24.69 39.05 63.93 37.05 51.06 60.09 31.62 18.55 47.02 64.63 38.93 41.07 37.35 45.56 36.95
1.25 23.26 38.99 63.13 41.87 43.79 59.06 33.76 20.66 65.93 100.8 45.44 43.09 50.5 51.02 51.02
1.5 24.77 42 79.1 64.98 96.36 64.11 33.47 21.65 89.63 103.76 59.41 49.25 76.25 69.79 78.28
1.75 24.54 39.08 71.37 74.64 98.74 59.68 31.85 19.3 108.56 158.95 70.25 52.11 95.57 103.41 101.76
2 23.43 39.83 66.41 47.75 89.98 61.15 34.14 20.65 116.68 208.11 74.49 49.51 96.45 79.35 93.46
0 20.74 38.4 57.65 36.75 40.48 56.42 24.32 24.65 7.55 7.69 24.02 24.18 11.67 20.97 11.58
0.25 17.9 34.03 52.44 32.03 40.27 48.91 23.74 22.99 10.09 21.26 26.64 27.86 12.82 21.18 14.01
0.5 18.72 33.18 51.65 27.34 39 47.32 19.87 16.82 16.62 58.54 29.97 33.15 17.48 26.29 19.55
0.75 19.19 34.73 56.67 34.18 43.26 49.95 25.81 17.5 29.86 103.45 35.11 38.22 28.17 31.97 30.87
3500 1 21.22 38.5 65.98 39.1 76.1 56.8 31.11 18.87 49.48 71.97 40.2 42.44 39.8 39.28 40.99
1.25 20.38 36.03 53.88 33.08 49.66 51.35 33.33 17.39 70.98 111.6 47.92 43.2 57.07 62.41 51.49
1.5 19.63 35.95 70.68 58.7 82.49 51.67 30.39 18.97 88.1 120.91 55.55 45.01 66.91 66.19 71.98
1.75 19.63 35.55 54.55 41.46 86.55 50.35 31.43 15.85 103.74 102.29 63.82 46.56 81.41 76.27 78.81
2 20.44 37.34 52.04 33.9 70.82 53.78 34.3 16.56 111.67 136.3 72.53 47.06 97.04 80.99 78.24
0 18.16 33.76 59.42 31.65 39.66 47.26 20.71 19.68 7.6 7.74 23.72 23.6 11.52 19.88 11.61
0.25 18.81 34.99 49.79 31.67 36.22 49.2 19.26 19.72 9.88 20.01 27.91 28.19 12.14 21.69 13.87
0.5 18.02 33.88 54.92 35.51 39.12 49.27 21.67 16.76 16.74 87.34 31.02 34.38 18.56 24.05 21.31
0.75 17.66 33.95 44.95 25.29 38.93 47.43 23.59 16.58 30.2 89.63 36.12 39.9 27.66 44.06 30.68
5000 1 17.48 33.63 49.97 29.1 48.54 47.89 28.7 18.13 46.43 80.45 40.07 42.06 39.98 46.77 38.25
1.25 20.04 36.54 53.89 33.91 50.57 51.67 32.42 16.86 72.71 121.44 49.42 45.18 58.9 66.74 56.3
1.5 17.68 32.9 45.59 30.67 60.29 45.23 31.14 14.83 92.41 162.72 56.67 43.41 74.65 78.07 63.22
1.75 16.91 32.38 47.65 36.18 80.13 45.23 27.74 16.96 105.55 123.41 66.73 48.54 86.49 92.77 83.16
2 17.7 33.02 48.57 36.72 104.62 47.78 30.67 14.75 117.25 147.32 71.05 47.19 94.93 91.72 96.34
@@
Train Shift CForest S-learner DR-learner T-learner CForest S-learner DR-l T-learner
size degree OS OS OS Ridge Tree OS Ridge Tree RCT wRCT RCT wRCT RCT RCT wRCT
0 -5.53 -6.38 -7.68 -6.73 -6.81 -9.25 -5.43 -5.59 0.03 0 0.07 0.05 -0.05 0.12 0.11
0.25 -5.37 -6.13 -7.43 -4.82 -5.46 -9.03 -5.33 -5.31 0.7 -0.78 0.75 0.24 0.47 0.84 0.31
0.5 -5.51 -6.22 -7.39 -5.89 -5.74 -9.2 -5.3 -4.66 1.73 -1.71 1.55 0.6 1.61 1.95 0.93
0.75 -5.71 -6.56 -7.68 -7.19 -6.15 -9.6 -6.08 -4.79 3.04 -1.09 2.29 0.77 2.01 2.95 1.58
500 1 -5.72 -6.46 -7.19 -5.86 -4.86 -9.36 -6.08 -4.18 4.73 -2.66 3.17 0.83 2.46 4.13 2.06
1.25 -5.36 -5.96 -6.53 -5.11 -4.3 -8.72 -5.8 -3.31 6.21 -2.69 4.3 1.05 4.06 5.46 3.24
1.5 -5.55 -6.2 -5.96 -4.16 -3.03 -9.24 -6.04 -3.22 7.82 2.47 5.49 1.68 4.29 6.67 4.99
1.75 -5.46 -6.35 -7.05 -5.34 -4.84 -9.26 -6.12 -2.65 9.31 6.51 6.98 2.38 6.17 8.49 7.01
2 -5.61 -6.45 -7.51 -5.71 -3.08 -9.38 -6.11 -2.29 9.97 4.22 7.23 2.79 6.33 8.59 6.87
0 -4.79 -6.39 -5.9 -3.77 -3.07 -8.12 -5.6 -6.46 -0.02 -0.05 -0.01 -0.05 0 -0.05 -0.06
0.25 -4.45 -5.98 -5.77 -2.93 -3.04 -7.63 -4.14 -4.08 0.64 -1.15 0.66 0.12 0.57 0.72 0.11
0.5 -4.8 -6.45 -5.86 -4.22 -3.9 -8.1 -4.8 -4.22 1.76 -2.54 1.46 0.45 2.06 1.8 0.64
0.75 -4.6 -6.2 -5.22 -2.59 -3 -7.92 -5.15 -4.06 3.27 -3.66 2.42 0.62 2.39 3.3 1.01
2000 1 -4.77 -6.39 -6.19 -4.61 -2.14 -8.03 -5.67 -3.56 4.77 0.86 3.11 0.6 2.93 4.18 2.43
1.25 -4.68 -6.29 -5.66 -3.2 -1.78 -8.09 -5.92 -3.12 6.73 1.77 4.66 1.22 3.23 6.07 4.55
1.5 -4.77 -6.3 -5.94 -2.86 -1.87 -7.96 -5.99 -2.81 7.87 4.02 5.75 1.41 4.91 7 5.68
1.75 -4.7 -6.32 -5.24 -3 -0.76 -8.07 -5.82 -2.3 9.27 3.31 6.61 1.9 5.11 8.01 6.33
2 -4.65 -6.32 -5.44 -3.35 -2.33 -8 -5.75 -1.88 10.11 0.68 7.17 2.21 8.35 8.59 4.95
0 -4.32 -5.94 -4.98 -2.43 -2.51 -7.37 -4.13 -4.27 0.03 -0.03 -0.03 -0.04 -0.08 0.03 0.01
0.25 -4.45 -6.06 -5.05 -2.79 -2.07 -7.54 -4.71 -4.5 0.71 -1.46 0.73 0.07 0.83 0.74 0.05
0.5 -4.22 -5.83 -4.95 -2.34 -1.98 -7.25 -4.33 -3.69 1.84 -2.67 1.43 0.3 1.72 1.82 0.65
0.75 -4.46 -6.16 -5.13 -2.75 -1.54 -7.6 -5.08 -3.64 3.15 -2.28 2.39 0.64 2.52 3.14 1.68
3500 1 -4.66 -6.31 -5.5 -2.87 -1.3 -7.72 -5.62 -3.48 4.92 -2.06 3.43 0.67 2.96 4.55 2.02
1.25 -4.26 -5.88 -4.72 -2.18 0.16 -7.29 -5.45 -2.86 6.58 -1.34 4.34 0.95 4.54 5.46 3.22
1.5 -4.2 -5.89 -4.68 -1.8 -0.39 -7.27 -5.42 -2.37 7.97 1.4 5.71 1.23 5.08 7.06 4.03
1.75 -4.48 -6.26 -5.09 -2.33 -0.46 -7.63 -5.78 -2.3 9.11 4.94 6.44 1.7 4.99 7.78 6.21
2 -4.47 -6.27 -5.68 -1.73 -1.48 -7.71 -5.75 -1.6 10.05 2.55 7.11 2.69 6.69 8.55 6.67
0 -4.28 -5.94 -4.76 -2.6 -2.26 -7.22 -3.78 -4.05 0.07 0.05 0 0.01 -0.11 -0.04 -0.05
0.25 -4.14 -5.88 -4.72 -2.49 -2.18 -7.1 -3.99 -3.83 0.72 -1.42 0.73 0.05 1.2 0.84 0.19
0.5 -4.13 -5.76 -4.92 -2.25 -1.93 -7.03 -4.24 -3.65 1.71 -2.44 1.44 0.26 2.13 1.85 0.46
0.75 -4.3 -5.96 -5.1 -2.23 -1.5 -7.17 -4.75 -3.52 3.17 -0.9 2.21 0.47 2.19 2.92 1.27
5000 1 -4.27 -5.89 -4.97 -1.93 -1.46 -7.1 -5.18 -3.11 4.89 -0.22 3.31 0.7 2.68 4.36 1.89
1.25 -4 -5.64 -4.34 -2.04 0.12 -6.86 -5.33 -2.63 6.54 4.26 4.45 0.92 4.74 5.65 4.17
1.5 -4.3 -6.05 -5.1 -2.42 -0.52 -7.28 -5.74 -2.49 8.24 3.39 5.82 1.43 4.18 7.26 5.58
1.75 -4.52 -6.16 -5.04 -2.59 -1.04 -7.4 -5.88 -2.11 9.14 4.9 6.47 1.98 5.79 7.9 6.7
2 -4.27 -6.02 -4.84 -1.91 -0.56 -7.23 -5.62 -1.67 9.96 3.5 7 1.97 6.6 8.45 6.43

### v2

@@ Results
Train Shift KL CForest S-learner DR-learner T-learner CForest S-learner DR-l T-learner
size degree Div. OS OS OS Ridge Tree OS Ridge Tree RCT wRCT RCT wRCT RCT RCT wRCT
0 2.82 37.43 47.2 94.84 100.75 76.41 86.4 30.54 33.85 7.71 7.76 24.31 24.37 11.11 21.63 11.12
0.25 2.94 37.22 47.11 95.35 75.57 81.38 87.48 31.09 31.7 9.33 11.57 26.49 26.4 11.53 19.14 11.52
0.5 3.30 38.09 47.77 93.25 70.52 66.67 84.78 35.26 35.56 17.13 34.55 33.46 35.15 18 24.62 17.87
0.75 4.07 35.07 44.28 93.6 91.21 64.14 78.57 30.76 29.05 27.5 52.74 34.61 37.01 25.26 32.78 26.67
500 1 5.44 38.69 49.32 102.71 95.48 166.78 91.69 37.94 31.16 47.8 84.65 41.38 43.62 40.63 40.01 43.95
1.25 6.95 34.34 43.26 92.51 77.29 69.69 81.17 36.15 26.93 67.26 158.72 47.33 42.52 55.63 49.43 52.96
1.5 9.37 34.78 43.29 117.41 64.09 82.92 84.48 34.91 24.9 88.28 141.6 54.21 43.13 65.94 70.72 55.18
1.75 12.16 36.48 44.88 86.79 84.26 77.4 77.93 36.47 25.58 104.41 134.25 66.15 46.86 85.97 89.8 78.45
2 15.04 40.22 51.41 117.57 103.47 108.49 91.47 35.91 27.33 117.4 166.51 73.7 51.99 96.27 91.76 85.72
0 2.52 24.77 39.92 69.39 38.64 44.04 60.21 28.88 27.3 8.2 8.49 25.29 25.55 11.9 20.17 12.31
0.25 2.92 22.66 36.62 61.15 46.68 40.69 56.5 21.7 21.79 9.11 17.4 25.27 25.67 12.27 19.06 13.67
0.5 3.38 23.36 40.53 63.01 43.82 57.06 60.26 24.92 20.93 16.28 45.42 31.93 35.5 17.8 25.63 19.64
0.75 4.09 24.91 41.14 62.5 37.85 47.54 61.23 28.29 20.71 29.3 72.08 35.42 39.42 27.35 38.08 30.23
2000 1 5.26 24.69 39.05 63.93 37.05 51.06 60.09 31.62 18.55 47.02 64.63 38.93 41.07 37.35 45.56 36.95
1.25 6.98 23.26 38.99 63.13 41.87 43.79 59.06 33.76 20.66 65.93 100.8 45.44 43.09 50.5 51.02 51.02
1.5 9.43 24.77 42 79.1 64.98 96.36 64.11 33.47 21.65 89.63 103.76 59.41 49.25 76.25 69.79 78.28
1.75 11.93 24.54 39.08 71.37 74.64 98.74 59.68 31.85 19.3 108.56 158.95 70.25 52.11 95.57 103.41 101.76
2 15.22 23.43 39.83 66.41 47.75 89.98 61.15 34.14 20.65 116.68 208.11 74.49 49.51 96.45 79.35 93.46
0 2.93 20.74 38.4 57.65 36.75 40.48 56.42 24.32 24.65 7.55 7.69 24.02 24.18 11.67 20.97 11.58
0.25 2.85 17.9 34.03 52.44 32.03 40.27 48.91 23.74 22.99 10.09 21.26 26.64 27.86 12.82 21.18 14.01
0.5 3.39 18.72 33.18 51.65 27.34 39 47.32 19.87 16.82 16.62 58.54 29.97 33.15 17.48 26.29 19.55
0.75 4.01 19.19 34.73 56.67 34.18 43.26 49.95 25.81 17.5 29.86 103.45 35.11 38.22 28.17 31.97 30.87
3500 1 5.27 21.22 38.5 65.98 39.1 76.1 56.8 31.11 18.87 49.48 71.97 40.2 42.44 39.8 39.28 40.99
1.25 6.79 20.38 36.03 53.88 33.08 49.66 51.35 33.33 17.39 70.98 111.6 47.92 43.2 57.07 62.41 51.49
1.5 9.18 19.63 35.95 70.68 58.7 82.49 51.67 30.39 18.97 88.1 120.91 55.55 45.01 66.91 66.19 71.98
1.75 11.94 19.63 35.55 54.55 41.46 86.55 50.35 31.43 15.85 103.74 102.29 63.82 46.56 81.41 76.27 78.81
2 15.15 20.44 37.34 52.04 33.9 70.82 53.78 34.3 16.56 111.67 136.3 72.53 47.06 97.04 80.99 78.24
0 2.64 18.16 33.76 59.42 31.65 39.66 47.26 20.71 19.68 7.6 7.74 23.72 23.6 11.52 19.88 11.61
0.25 2.92 18.81 34.99 49.79 31.67 36.22 49.2 19.26 19.72 9.88 20.01 27.91 28.19 12.14 21.69 13.87
0.5 3.33 18.02 33.88 54.92 35.51 39.12 49.27 21.67 16.76 16.74 87.34 31.02 34.38 18.56 24.05 21.31
0.75 4.18 17.66 33.95 44.95 25.29 38.93 47.43 23.59 16.58 30.2 89.63 36.12 39.9 27.66 44.06 30.68
5000 1 5.23 17.48 33.63 49.97 29.1 48.54 47.89 28.7 18.13 46.43 80.45 40.07 42.06 39.98 46.77 38.25
1.25 6.87 20.04 36.54 53.89 33.91 50.57 51.67 32.42 16.86 72.71 121.44 49.42 45.18 58.9 66.74 56.3
1.5 9.18 17.68 32.9 45.59 30.67 60.29 45.23 31.14 14.83 92.41 162.72 56.67 43.41 74.65 78.07 63.22
1.75 12.10 16.91 32.38 47.65 36.18 80.13 45.23 27.74 16.96 105.55 123.41 66.73 48.54 86.49 92.77 83.16
2 15.23 17.7 33.02 48.57 36.72 104.62 47.78 30.67 14.75 117.25 147.32 71.05 47.19 94.93 91.72 96.34
@@
Train Shift KL CForest S-learner DR-learner T-learner CForest S-learner DR-l T-learner
size degree Div. OS OS OS Ridge Tree OS Ridge Tree RCT wRCT RCT wRCT RCT RCT wRCT
0 2.98 -5.53 -6.38 -7.68 -6.73 -6.81 -9.25 -5.43 -5.59 0.03 0 0.07 0.05 -0.05 0.12 0.11
0.25 3.19 -5.37 -6.13 -7.43 -4.82 -5.46 -9.03 -5.33 -5.31 0.7 -0.78 0.75 0.24 0.47 0.84 0.31
0.5 3.62 -5.51 -6.22 -7.39 -5.89 -5.74 -9.2 -5.3 -4.66 1.73 -1.71 1.55 0.6 1.61 1.95 0.93
0.75 4.57 -5.71 -6.56 -7.68 -7.19 -6.15 -9.6 -6.08 -4.79 3.04 -1.09 2.29 0.77 2.01 2.95 1.58
500 1 5.78 -5.72 -6.46 -7.19 -5.86 -4.86 -9.36 -6.08 -4.18 4.73 -2.66 3.17 0.83 2.46 4.13 2.06
1.25 7.79 -5.36 -5.96 -6.53 -5.11 -4.3 -8.72 -5.8 -3.31 6.21 -2.69 4.3 1.05 4.06 5.46 3.24
1.5 10.33 -5.55 -6.2 -5.96 -4.16 -3.03 -9.24 -6.04 -3.22 7.82 2.47 5.49 1.68 4.29 6.67 4.99
1.75 13.82 -5.46 -6.35 -7.05 -5.34 -4.84 -9.26 -6.12 -2.65 9.31 6.51 6.98 2.38 6.17 8.49 7.01
2 16.98 -5.61 -6.45 -7.51 -5.71 -3.08 -9.38 -6.11 -2.29 9.97 4.22 7.23 2.79 6.33 8.59 6.87
0 2.99 -4.79 -6.39 -5.9 -3.77 -3.07 -8.12 -5.6 -6.46 -0.02 -0.05 -0.01 -0.05 0 -0.05 -0.06
0.25 3,00 -4.45 -5.98 -5.77 -2.93 -3.04 -7.63 -4.14 -4.08 0.64 -1.15 0.66 0.12 0.57 0.72 0.11
0.5 3.67 -4.8 -6.45 -5.86 -4.22 -3.9 -8.1 -4.8 -4.22 1.76 -2.54 1.46 0.45 2.06 1.8 0.64
0.75 4.61 -4.6 -6.2 -5.22 -2.59 -3 -7.92 -5.15 -4.06 3.27 -3.66 2.42 0.62 2.39 3.3 1.01
2000 1 5.99 -4.77 -6.39 -6.19 -4.61 -2.14 -8.03 -5.67 -3.56 4.77 0.86 3.11 0.6 2.93 4.18 2.43
1.25 7.48 -4.68 -6.29 -5.66 -3.2 -1.78 -8.09 -5.92 -3.12 6.73 1.77 4.66 1.22 3.23 6.07 4.55
1.5 10.33 -4.77 -6.3 -5.94 -2.86 -1.87 -7.96 -5.99 -2.81 7.87 4.02 5.75 1.41 4.91 7 5.68
1.75 13.73 -4.7 -6.32 -5.24 -3 -0.76 -8.07 -5.82 -2.3 9.27 3.31 6.61 1.9 5.11 8.01 6.33
2 17.68 -4.65 -6.32 -5.44 -3.35 -2.33 -8 -5.75 -1.88 10.11 0.68 7.17 2.21 8.35 8.59 4.95
0 2.99 -4.32 -5.94 -4.98 -2.43 -2.51 -7.37 -4.13 -4.27 0.03 -0.03 -0.03 -0.04 -0.08 0.03 0.01
0.25 2.97 -4.45 -6.06 -5.05 -2.79 -2.07 -7.54 -4.71 -4.5 0.71 -1.46 0.73 0.07 0.83 0.74 0.05
0.5 3.50 -4.22 -5.83 -4.95 -2.34 -1.98 -7.25 -4.33 -3.69 1.84 -2.67 1.43 0.3 1.72 1.82 0.65
0.75 4.45 -4.46 -6.16 -5.13 -2.75 -1.54 -7.6 -5.08 -3.64 3.15 -2.28 2.39 0.64 2.52 3.14 1.68
3500 1 6.00 -4.66 -6.31 -5.5 -2.87 -1.3 -7.72 -5.62 -3.48 4.92 -2.06 3.43 0.67 2.96 4.55 2.02
1.25 7.50 -4.26 -5.88 -4.72 -2.18 0.16 -7.29 -5.45 -2.86 6.58 -1.34 4.34 0.95 4.54 5.46 3.22
1.5 10.40 -4.2 -5.89 -4.68 -1.8 -0.39 -7.27 -5.42 -2.37 7.97 1.4 5.71 1.23 5.08 7.06 4.03
1.75 13.87 -4.48 -6.26 -5.09 -2.33 -0.46 -7.63 -5.78 -2.3 9.11 4.94 6.44 1.7 4.99 7.78 6.21
2 17.16 -4.47 -6.27 -5.68 -1.73 -1.48 -7.71 -5.75 -1.6 10.05 2.55 7.11 2.69 6.69 8.55 6.67
0 3.12 -4.28 -5.94 -4.76 -2.6 -2.26 -7.22 -3.78 -4.05 0.07 0.05 0 0.01 -0.11 -0.04 -0.05
0.25 3.06 -4.14 -5.88 -4.72 -2.49 -2.18 -7.1 -3.99 -3.83 0.72 -1.42 0.73 0.05 1.2 0.84 0.19
0.5 3.59 -4.13 -5.76 -4.92 -2.25 -1.93 -7.03 -4.24 -3.65 1.71 -2.44 1.44 0.26 2.13 1.85 0.46
0.75 4.56 -4.3 -5.96 -5.1 -2.23 -1.5 -7.17 -4.75 -3.52 3.17 -0.9 2.21 0.47 2.19 2.92 1.27
5000 1 5.82 -4.27 -5.89 -4.97 -1.93 -1.46 -7.1 -5.18 -3.11 4.89 -0.22 3.31 0.7 2.68 4.36 1.89
1.25 7.69 -4 -5.64 -4.34 -2.04 0.12 -6.86 -5.33 -2.63 6.54 4.26 4.45 0.92 4.74 5.65 4.17
1.5 10.31 -4.3 -6.05 -5.1 -2.42 -0.52 -7.28 -5.74 -2.49 8.24 3.39 5.82 1.43 4.18 7.26 5.58
1.75 13.72 -4.52 -6.16 -5.04 -2.59 -1.04 -7.4 -5.88 -2.11 9.14 4.9 6.47 1.98 5.79 7.9 6.7
2 17.32 -4.27 -6.02 -4.84 -1.91 -0.56 -7.23 -5.62 -1.67 9.96 3.5 7 1.97 6.6 8.45 6.43